### v2

# Cleaning WASP-33 b transits from the host star photometric variability: analysis of TESS data from two sectors

Abstract
Based on TESS observations of a $\delta$ Scuti variable WASP-33 obtained in 2019 and 2022, we thoroughly investigate the power spectrum of this target photometric flux and construct a statistically exhaustive model for its variability. This model contains $30$ robustly justified harmonics detected in the both TESS sectors simultaneously, $13$ less robust harmonics detected in a single TESS sector, a red noise and a quasiperiodic noise terms. This allowed us to greatly improve the accuracy of the exoplanet WASP-33 b transit timings, reducing the TTV residuals r.m.s. drastically, by a factor of $3.5$, from $63$ s to $18$ s. Finally, our analysis does not confirm existence of a detectable orbital phase variation claimed previously based on WASP-33 TESS photometry of 2019.
keywords:
techniques: photometric, stars: variables: Scuti, stars: individual: WASP-33, planets and satellites: detection
pagerange: Cleaning WASP-33 b transits from the host star photometric variability: analysis of TESS data from two sectorsReferencespubyear: 2024
1 Introduction
WASP-33 (HD 15082) is a bright 8.14 mag fast-rotating star with $v\sin i\sim 90$ km/s. Its only known exoplanet WASP-33 b was detected by SuperWASP (Herrero et al., 2011) with orbital period currently estimated by $P_{b}=(1.219867\pm 4.2\cdot 10^{-5})$ d (von Essen et al., 2020), mass by $M_{b}=(2.10\pm 0.14)$ $M_{\rm Jup}$, and radius by $r_{b}=(1.59\pm 0.07)$ $R_{\rm Jup}$ (Chakrabarty & Sengupta, 2019). It is remarkable that this exoplanet was the first one ever known to orbit a $\delta$ Scuti variable.
The variability of the prototype star itself, $\delta$ Scuti, was discovered in the beginning of the XX century (Campbell & Wright, 1900), later confirmed by spectral and photometric observations (Colacevich, 1935; Fath, 1935). Furthermore, Eggen (1956) combined this and a few similar stars together in a standalone class of variables, and so the $\delta$ Scuti type of variable stars appeared.
These variables are pulsating stars belonging to spectral classes A0-F5 III-IV, located in the instability strip of the HR diagram (Baglin et al., 1973; Breger & Stockenhuber, 1983). Their $m_{V}$ variation amplitudes range from $0.003$ to $0.9$ (typically a few hundreds) and periods from $0.01$ d to $0.2$ d. Most of the $\delta$ Scuti stars belong to type I population (Breger, 1979), with a typical mass of $\sim 2$ $M_{\odot}$ (Milligan & Carson, 1992). The power spectrum of $\delta$ Scuti variables may reveal tens of periods (Breger et al., 1999) ranging from hours to minutes. Such stars often exhibit significant changes in the form of their lightcurve, in their variability period and amplitude (Blazhko effect). A key role in the pulsation generating mechanism belongs to helium which is subject to heating/ionization and cooling/recombination processes, thus changing its opacity in a cycle (Cox, 1963). Radial as well as non-radial pulsations can be observed.
Large amount of data regarding $\delta$ Scuti variables were brought by (i) OGLE and MACHO surveys which detected about $15000$ such variables in the Large Magellanic Cloud (Poleski et al., 2010; Garg et al., 2010; Soszyński et al., 2023), (ii) Kepler/K2 missions (Borucki et al., 2010; Guzik, 2021) that found about $2000$ variables of this type, and (iii) TESS (Transiting Exoplanet Survey Satellite) mission, for example Gootkin et al. (2024) identified a sample of $\sim 14000$ variables likely belonging to the $\delta$ Scuti class.
The intrinsic variability of WASP-33 represents an obvious nuisance factor for analysing its exoplanet b transit lightcurves. After the discovery work by Herrero et al. (2011), the effect of this variability on the estimated planetary orbital parameters was investigated using out-of-transit observations from ground-based telescopes located in Germany and Spain (von Essen et al., 2014). Their investigation of the WASP-33 pulsations power spectrum revealed eight significant frequencies. It was concluded that cleaning the transit lightcurves from stellar pulsations practically does not affect estimated planetary parameters, but their uncertainties get decreased. Besides, von Essen et al. (2014) noticed that WASP-33 pulsations phases change over time. In a later analysis of TESS photometry of 2019, von Essen et al. (2020) claimed the detection of $29$ pulsation frequencies. In addition to the power spectrum analysis, they also reported statistically significant detection of WASP-33 b secondary eclipse and of the orbital phase curve variation (period of $P_{b}$, amplitude of $100.4\pm 13.1$ ppm). Also, they took into account the “ellipsoidal variation” that appears because of planetary gravitational effect on the star shape (period of $P_{b}/2$, theoretically predicted amplitude of $27.4$ ppm).
A more recent analysis of WASP-33 variability was performed using data from TESS (Kálmán et al., 2022). They investigated the Fourier spectrum of stellar oscillations reconstructed with TLCM (Transit and Light Curve Modeller, Csizmadia et al. 2023). This spectrum revealed three peaks with $S/N>8$ that were close to a commensurability with the planetary orbital period ($1:3$, $1:12$, and $1:25$). It was argued that the distribution of $17$ frequencies with the largest amplitudes tend to shift right from these subharmonics. As Kálmán et al. (2022) noticed, this effect can be caused by the tidal perturbations from the planet.
Our primary goal here is to analyse WASP-33 pulsations spectrum based on the up-to-date TESS data (observations of 2019 and 2022), and to investigate whether (and how much) this can help us to improve the transit timings accuracy for the exoplanet WASP-33 b. The paper contains three major sections, discussing subsequently the TESS observations in Sect. 2, construction of the photometric variation model in Sect. 3, and applying this model to improve the set of estimated planet parameters and transit timings, in Sect. 4.
2 TESS photometry of WASP-33
2.1 Basic setting
TESS observed WASP-33 in sector 18 (Nov 2019) and sector 58 (Nov 2022), lasting about $1$ month each. In the DVT data downloaded from the MAST (Mikulski Archive for Space Telescopes) we found $14796$ and $18951$ photometric measurements for these two sectors, respectively. We identified $17$ transits in sector 18 and $21$ transits in sector 58 (plus a partial transit in the sector end). For each transit we extracted a ±$\pm 4$ h piece of the lightcurve about the predicted midtransit epoch. This subset formed our “in-transit” dataset (ITD). The “out-of-transit” dataset (OTD) included all observations located at least ±$\pm 2$ h away from the centres of transits and secondary eclipses. Notice that ITD and OTD slightly overlap.
In this work we will deal mainly with the OTD, since we aim to construct a model for the intrinsic star variability. The sector 18 portion of the OTD will be hereafter refered to as TESS18, and it contains $10684$ datapoints. The sector 58 OTD portion (TESS58) contains $13625$ datapoints. ITD contains $8914$ datapoints in total, with $\sim 200$ observations per transit.
Figure 1: Spectral window of the TESS18 (left) and TESS58 (right) data, for three frequency ranges (from top to bottom). The (quasi-) Nyquist alias peak at $7202f_{\rm Ny}=$ 720 d-1 is seen in the bottom panels.
TESS data look close to being even with a constant timestep of $\Delta t=2$ min. However, this is not strictly fulfilled even within just a single sector, because in the middle of each sector there was a gap, after which the $2$ min grid pattern got shifted. Also, the OTD set contains holes in place of transits and secondary eclipses. But nevertheless the spectral window functions of TESS18,58 data suggest that these data can be treated as nearly even, in particular there is a classic aliasing pattern suggesting the existence of a “quasi-Nyquist“ frequency $f_{\rm Ny}=1/(2\Delta t)=360$ d-1 (see Fig. 1).
2.2 Choosing the TESS data product to process
A yet another important question is the choice of the photometric product for the analysis. For each target, MAST offers a rich line of photometric data that correspond to different levels of reduction and processing (Twicken et al., 2018; Li et al., 2019; Twicken et al., 2020; Jenkins, 2020). In Fig. 2 we plot all types of photometric data available for WASP-33. von Essen et al. (2020) considered two lowest-level types of data, namely SAP (Simple Aperture Photometry) and PDCSAP (Presearch Data Conditioning SAP), and selected the raw SAP one motivating this choice by the increased level of correlated noise revealed in PDCSAP. They estimated this increase by a factor of $\sim 2.4$ using the $\beta$ method by Winn et al. (2008) (their citation to Carter & Winn 2009 appears incorrect). It is simultaneously recognised that WASP-33 stellar pulsations cause larger photometric effect than possible PDC-generated correlated noise, and in such a case so large increase factor as $2.4$ looks relatively puzzling. However, the $\beta$ method they used deals with relative noise contributions. The relative portion of a non-white noise (NWN hereafter) may indeed increase whenever PDC added some, but alternatively this same picture would appear if PDC removed a portion of the white noise (WN hereafter), keeping the NWN part intact.
Figure 2: Different types of TESS photometric data products of WASP-33, for the Sector 18 (left) and Sector 58 (right); the data types are given in the ordinate labels. Top four panels additionally show a trend curve (black) obtained by smoothing the OTD with a Gaussian kernel (sigma of $1$ d). In these four plots we also print the r.m.s. of the OTD flux divided by the smooth model. All DVT graphs show the r.m.s. of the corresponding OTD (which were already normalized) too. The r.m.s. are comparable between each other from top to the fourth row, while in the bottom plots they are close to unity by whitening.
Indeed, let us look at the raw SAP data (Fig. 2, top panels). They reveal obvious large-scale trend which is removed by PDC (second pair of panels). However, the PDC effect is not limited to only detrending. To verify this, we approximated the trend by smoothing the SAP OTD with a Gaussian kernel (width of $1$ d), then divided it by this simple model, and computed the r.m.s. of the resulting normalized flux. For the sector 18 data this resulted in $1.73$ parts per thousand (ppt). Applying the same procedure to the PDCSAP OTD we obtained the r.m.s. of $1.38$ ppt. One can see that additionally to the trend, the PDC removed $\sim 1.04$ ppt in terms of standard deviation (only in sector 18 data; in sector 58 the r.m.s. appeared nearly identical). Therefore, based on this preliminary analysis we believe that discarding PDCSAP in favour of the raw SAP data does not look justified enough, in this particular case at least. Even if PDC did add some extra noise to the data, it simultaneously removed more.
Beyond the PDC, the DVT (Data Validation Timeseries) are publicly released. They involve three levels of reduction, which can be refered to by the corresponding column titles in the DVT files: LC_INIT, LC_WHITE, and LC_DETREND. Briefly, the DVT involve (i) optional harmonics removal, (ii) edge detrending, (iii) level adjustment, (iv) normalization, (v) gap unfilling, and (vi) stitching between sectors (Twicken et al., 2020). The LC_INIT data for WASP-33 (see third pair of plots in Fig. 2) have only a minimal difference from the PDCSAP. This can be verified by considering the graphs of the ratio ${\rm PDCSAP}/(1+{\rm LC\_INIT})$, which are shown in the second pair of plots in Fig. 2. We can see that LC_INIT corrects for minor edge-jump effects that are likely tied to the satellite momentum dumps every few days, which were also synchronized with pauses made in observations for data download. [^1] See Data Release Notes at https://archive.stsci.edu/missions/tess/doc/tess_drn/tess_sector_18_drn25_v02.pdf and https://archive.stsci.edu/missions/tess/doc/tess_drn/tess_sector_58_drn83_v02.pdf We believe that this effect is important for our study and should be necessarily removed. To be honest, its reduction provided with the LC_INIT data is not always perfect, see e.g. the end of sector 58 (Fig. 2, right, second panel) that was fit badly because of a partial transit. However, building a more efficient model for this effect would be out of the scope of current work. That problematic partial transit was removed from our OTD anyway, and it was not included in the ITD data as well, so the corrupted portion of data did not affect our analysis.
Concerning the columns LC_WHITE and LC_DETREND, they involve more substantial changes to the flux. It seems these reductions perform at least qualitatively similar to our analysis, so they will interfere with it, so we selected the DVT LC_INIT data.
The final notice is that the “third light” correction is already applied at the PDC step, so we do not need to apply it here. There are three known sources that jointly provide about $2$ per cent extra flux inside the photometric aperture of WASP-33 (von Essen et al., 2020). Their estimation agreed with the TESS-provided one within $\sim 10$ per cent (the CROWDSAP field stored in the FITS file).
3 Building the statistically exhaustive model of WASP-33 variability
3.1 Models
Here our goal is to construct a statistically “exhaustive” WASP-33 model that represents its OTD so that the residuals do not contain statistically significant periodic (or quasiperiodic) variations. In this attempt we aimed to apply a periodogram-based approach somewhat reminiscent of the CLEANest algorithm by Foster (1995). Namely, we represent the photometric variation as a sum of sinusoidal harmonics:
$\mu_{n}(t)=A_{0}+\sum_{k=1}^{n}A_{k}\cos\left(2\pi f_{k}(t-T_{0})+\varphi_{k}\right),$ (1)
where $A_{k}$, $\varphi_{k}$, and $f_{k}=1/P_{k}$ are free parameters.
However, if the model (1) was used in the white noise treatment then we would likely need hundreds of harmonics to represent the OTD exhaustively. After extraction of some initial harmonics ($5$ to $7$ of the most significant ones), the residual periodogram did not reveal a single clearly dominating peak, looking more similar to a spectrum of banded NWN.
Alone such an observation is not yet decisive to claim that the data physically involve NWN. Two alternative models “a lot of harmonics + WN” and “fewer harmonics + NWN” may appear formally indistinguishable, i.e. statistically equivalent in how well they describe the data. But even in this case the NWN model may offer a great reduction in the number of model components (and parameters), so we can profit from a mathematically more simple $\mu$.
To take the NWN into account, we adopted an approach of Gaussian process (GP) fitting, introduced in (Baluev, 2011). In this approach we should define a model for the NWN correlation function, which we now set to:
$k_{n^{\prime}}(\Delta t_{ij})$ $=\left(\sigma_{{\rm TESS},i}^{2}+\sigma_{0}^{2}\right)\delta_{ij}+\sum_{k=1}^{n^{\prime}}\sigma_{k}^{2}{\rm e}^{-\frac{|\Delta t_{ij}|}{\tau_{k}}}\cos 2\pi f^{\prime}_{k}\Delta t_{ij},$
$\phantom{=}\Delta t_{ij}=t_{i}-t_{j},$ (2)
where $\sigma_{{\rm TESS},i}$ is TESS-provided photometric uncertainty at $t_{i}$, $\delta_{ij}$ is Kronecker delta, while $\sigma_{k}$, $\tau_{k}$, and $f^{\prime}_{k}=1/P^{\prime}_{k}$ are free parameters. This model involves the WN portion ($\sigma_{0}$) and $n^{\prime}$ independent quasiperiodic components (purely NWN part). By a visual look at the periodograms (e.g. see Fig. 3 discussed below) we noticed three possible spectrum bands that are concentrated near the frequencies $f^{\prime}_{1}\simeq 0$, $f^{\prime}_{2}\simeq 10$ d-1, and $f^{\prime}_{3}\simeq 20$ d-1. We set these frequencies as starting values for the GP fitting, thus assuming $n^{\prime}=3$.
The models (1) and (2) can be fitted jointly using the maximum likelihood approach, assuming that noise is a Gaussian random process (Baluev, 2011, 2013b). It is based on minimizing the function
$f(\mn@boldsymbol{\theta},\mn@boldsymbol{\eta})=\log\det\textbf{{V}}(\mn@boldsymbol{\eta})+\frac{1}{\gamma}\,{\mn@boldsymbol{r}}^{\mathrm{T}}(\mn@boldsymbol{\theta})\,\textbf{{V}}^{-1}(\mn@boldsymbol{\eta})\,\mn@boldsymbol{r}(\mn@boldsymbol{\theta})\;\stackrel{{\scriptstyle\mn@boldsymbol{\theta},\mn@boldsymbol{\eta}}}{{\longmapsto}}\;\min,$
${r}=\mn@boldsymbol{x}-\mn@boldsymbol{\mu}(\mn@boldsymbol{t},\mn@boldsymbol{\theta}),\quad\gamma=1-\frac{\dim\mn@boldsymbol{\theta}}{\dim\mn@boldsymbol{x}},$ (3)
where we have (i) input data including observation times ${t}$ and the flux ${x}$, (ii) the models ${\mu}$ being one of (1) and V which elements are filled in with (2), and (iii) vectors of fittable parameters ${\theta}$ (which includes $A_{k}$, $\varphi_{k}$, and $f_{k}$) and ${\eta}$ (which includes $\sigma_{k}^{2}$, $\tau_{k}$, and $f_{k}^{\prime}$). The correction factor $\gamma$ is necessary to reduce the bias in ${\eta}$.
3.2 Algorithm
We run a sequential periodogram algorithm according to the following scheme. We started from $k_{3}$ as specified above, and from $\mu_{2}$ involving two harmonics with fixed frequencies as per von Essen et al. (2020): $f_{1}=P_{b}$ (orbital phase variation) and $f_{2}=P_{b}/2$ (ellipsoidal variation). After that the model (1) can be built up iteratively by adding harmonics that refer to the tallest peak in the subsequent residual periodogram. These periodograms must take the NWN into account via (2). For that we used the formalism of the likelihood-ratio periodogram (LRP) introduced by Baluev (2009, 2013b) and implemented in the PlanetPack software (Baluev, 2013a, 2018). The LRP considers two rival models of the data, a base one ($\mu_{n}$ on $n$th iteration) and an alternative one involving an extra sinusoid (i.e. $\mu_{n+1}$). The likelihood ratio statistic based on (3) is then computed for these models and its logarithm is plotted against the frequency of the last added sinusoid. Such a graph represents our LRP that generalises the classic Lomb (1976) – Scargle (1982) periodogram. The LRP can be computed for the classic WN case as well as for the NWN model (2), by defining the likelihood function (namely, the model $\textbf{{V}}(\mn@boldsymbol{\eta})$) appropriately. After each iteration all free parameters of so far detected harmonics are refined by refitting with the new model $\mu$.
Very importantly, the spectral window in Fig. 1 does not demonstrate large side peaks (in the Nyquist range at least). The height of the tallest side peak is $\simeq 0.1$. Therefore, the aliasing appears insignificant in our task, and we may safely assume that the maximum periodogram peak corresponds to the true (rather than alias) frequency. This makes all the analysis easier, for example it is unnecessary to care about the order in which the harmonics are extracted (which could be important otherwise).
However, dealing with NWN periodograms implies numerous NWN model refitting, which is computationally demanding. The likelihood function involves operations on very large covariance matrices ($\sim 10^{4}\times 10^{4}$), rendering the LRP computation too slow for being accomplished in full. But we tried to bypass this limitation using the following approach. On each iteration we computed, at first, the fast WN-only version of the LRP, and determined its $256$ tallest peaks. Then for each of these peaks we run an NWN maximum-likelihood fitting, starting it from the associated WN best fit. This allowed us to compute heights of these same peaks in the NWN treatment, and then to select the tallest one among them.
Such an approach is potentially vulnerable because it may, in theory, miss the actual best-significance harmonic. Even $256$ peaks that appeared tallest with the WN model do not necessarily contain the one that should appear tallest in the NWN case. However, the probability of such a condition decreases when the number of peaks being inspected grows, so the validity of the results can be tested by a probe increase of that number (in the end of iterations sequence).
We launched our harmonics detection iterations assuming the frequency range from $f=0$ to $f=50$ d-1, which is a priori reasonable for $\delta$ Scuti stars, and also contained all apparently noticeable periodogram power. As we specified above, the NWN model (2) was initially set to have $n^{\prime}=3$ terms, but somewhere in the middle of the sequence the third NWN term (with $f^{\prime}\simeq 20$ d-1) appeared, basically, decomposed into harmonics. We started to obtain various degeneracy issues regarding this term, for example it was trying to either model just a single sinusoid ($\tau\to 0$), or to mimic a portion of the white noise ($\tau\to\infty$), or to model a portion of another NWN term. When this behaviour started, we manually removed this third term from the NWN model, and continued the sequence further with $k_{2}(\Delta t)$ instead of $k_{3}(\Delta t)$. Finally, when no significant harmonics left (in the NWN treatment), we reprocessed the last iteration inspecting $1024$ tallest peaks instead of just $256$, and in the entire Nyquist range (max frequency of $360$ d-1 instead of $50$ d-1). This did not result in detection of additional harmonics, ensuring that we did not miss any.
Figure 3: Periodogram analysis and extraction of the harmonics. The full set of figures is available in the online-only supplement.
This algorithm was applied to TESS18 and TESS58 data separately, which resulted in quite similar iterations sequences and in similar results. Some graphs illustrating the initial and final iterations are shown in Fig. 3. Each panel demonstrates the WN periodograms (raw as well as smoothed one), the theoretic power spectrum of the current best fitting NWN model based on (2), and the set of NWN peaks that were dealt with on this iteration. On the last iteration, when we inspected $1024$ peaks, those peaks are covering the entire Nyquist range without significant holes or blind segments, so it is very unlikely that some statistically significant harmonics escaped our analysis. Please also notice remarkable similarity between our fitted NWN power spectrum and the smoothed WN periodogram.
This algorithm belongs to a frequentist type, and is based, in its core, on the likelihood ratio test. We did not try methods of a Bayesian type here, because it was shown by Baluev (2022) that in the task of sinusoidal signal detection, without significant aliasing at least, detection efficiency of a Bayesian analysis (measured by certain formal metric) appears practically identical to the usual non-Bayesian periodograms.
3.3 Detection threshold
An explanation should be given regarding how we computed the statistical significance of periodogram peaks in the NWN treatment. As explained by Baluev (2009), for the WN type of the LRP we may use the following approximation of the False Alarm Probability (FAP) from (Baluev, 2008):
${\rm FAP}(z)\stackrel{{\scriptstyle z\gg 1}}{{\simeq}}W{\rm e}^{-z}\sqrt{z},\quad W=(f_{\max}-f_{\min})T_{\rm eff},$ (4)
where $z$ is the height of the tallest peak, and $T_{\rm eff}$ is an effective time span proportional to the sample variance of the timings $t_{i}$. As argued (without proofs) in (Baluev, 2015), the approximate formula (4) remains applicable for NWN models as well, but with $T_{\rm eff}$ somewhat differently determined. Delisle et al. (2020) developed this study further, partly confirming this conclusion and providing a more general definition for $T_{\rm eff}$. However, in our task it appeared that $T_{\rm eff}$ always remains practically identical, regardless of the noise model or which approximation to adopt. We obtained $T_{\rm eff}$ about $24$ d for TESS18 and about $28$ d for TESS58.
In this study we aimed to obtain a more accurate final model of the WASP-33 variability, possibly even by the cost of larger probability of false positive harmonics. So we adopted a relatively mild FAP threshold of $5\%$ (the iterations continued until the right hand side of (4) exceeded $0.05$).
3.4 Results
Table 1 lists all the harmonics that we detected in TESS18 and in TESS58.222Except for the two harmonics with fixed periods $P_{b}$ and $P_{b}/2$. Those are not shown in Table 1 for clarity, and are discussed separately. It is highly important that both these sectors revealed very similar sets of the harmonics. Nearly all harmonics can be cross-identified between TESS18 and TESS58 by comparing their frequencies (honouring uncertainties). Usually, the amplitudes of the cross-identified harmonics appeared similar as well, although some of $A_{k}$ demonstrated big changes between TESS18 and TESS58. It is impossible to conclusively compare the phases of these harmonics, because of too large uncertainty accumulated over three years passed between these TESS sectors. In general, one may conclude that WASP-33 variability pattern demonstrates remarkable, though not entirely perfect, stability over the timescale of a few years at least.
Table 1: Estimated harmonics parameters.
TESS18 ($T_{0}=2458800$) TESS58 ($T_{0}=2459895$)
IDs $P$ [d] $A$ [ppm] $\varphi$ [] $P$ [d] $A$ [ppm] $\varphi$ []
$H_{1}$ $0.0213065(17)$ $41.7(8.0)$ $328(11)$
$H_{2}$ $F_{10}$ $0.02930411(56)$ $228.3(6.8)$ $297.9(1.8)$ $0.02930358(45)$ $312.5(8.3)$ $279.2(1.5)$
$H_{3}$ $F_{28}$ $0.0328267(44)$ $38.0(7.0)$ $306(11)$
$H_{4}$ $F_{26}$ $0.0331895(39)$ $44.5(7.1)$ $37.1(9.8)$ $0.0331986(43)$ $42.4(8.5)$ $155(12)$
$H_{5}$ $F_{22}$ $0.0348597(28)$ $69.6(7.2)$ $206.2(6.4)$ $0.0348553(16)$ $130.6(8.7)$ $186.8(3.8)$
$H_{6}$ $F_{8}$ $0.03597781(78)$ $267.3(7.3)$ $317.9(1.7)$ $0.03597797(38)$ $580.6(8.7)$ $93.27(87)$
$H_{7}$ $F_{18}$ $0.0364189(26)$ $84.0(7.3)$ $180.1(5.4)$ $0.0364150(26)$ $86.0(8.7)$ $315.9(5.9)$
$H_{8}$ $0.0373459(57)$ $42.3(8.8)$ $154(12)$
$H_{9}$ $F_{14}$ $0.0390044(23)$ $111.5(7.5)$ $198.8(4.2)$ $0.0390015(30)$ $87.9(8.9)$ $214.6(5.9)$
$H_{10}$ $F_{9}$ $0.0401869(11)$ $253.0(7.8)$ $11.5(1.9)$ $0.0401856(11)$ $258.3(9.1)$ $118.1(2.1)$
$H_{11}$ $0.0406911(67)$ $43.5(8.0)$ $9(11)$
$H_{12}$ $F_{16}$ $0.0430933(35)$ $102.1(8.2)$ $258.8(5.1)$ $0.0430910(30)$ $114.8(9.4)$ $55.1(4.8)$
$H_{13}$ $F_{27}$ $0.0440461(85)$ $41.3(8.7)$ $121(13)$ $0.0440554(71)$ $51(10)$ $221(11)$
$H_{14}$ $F_{20}$ $0.0451501(44)$ $90.3(8.5)$ $224.4(6.0)$ $0.0451491(44)$ $91.3(9.9)$ $245.1(6.1)$
$H_{15}$ $0.0454574(96)$ $40.6(8.6)$ $103(13)$
$H_{16}$ $0.0457087(92)$ $43.5(8.7)$ $60(12)$ $0.0457113(72)$ $56.2(9.8)$ $194(10)$
$H_{17}$ $F_{19}$ $0.0460159(54)$ $76.8(8.8)$ $261.9(7.0)$ $0.0460004(57)$ $72.1(9.8)$ $73.5(7.8)$
$H_{18}$ $0.0465026(83)$ $53.0(8.8)$ $296(10)$
$H_{19}$ $F_{24}$ $0.0469927(71)$ $61.5(8.8)$ $164.0(8.7)$ $0.0469969(49)$ $85.6(9.7)$ $268.1(6.5)$
$H_{20}$ $F_{4}$ $0.04747567(95)$ $617(11)$ $335.46(96)$ $0.04747626(78)$ $618(11)$ $161.2(1.0)$
$H_{21}$ $F_{11}$ $0.0476547(89)$ $167(22)$ $54.4(4.4)$ $0.0476614(36)$ $190(11)$ $42.4(3.4)$
$H_{22}$ $0.047747(12)$ $119(21)$ $86.7(5.9)$ $0.0477543(62)$ $102(11)$ $101.6(6.5)$
$H_{23}$ $F_{6}$ $0.0486902(15)$ $335.7(8.9)$ $259.1(1.7)$ $0.0486897(15)$ $330(10)$ $125.5(1.8)$
$H_{24}$ $F_{1}$ $0.04959774(65)$ $780.8(9.1)$ $1.13(72)$ $0.04959762(62)$ $786(10)$ $265.35(74)$
$H_{25}$ $F_{17}$ $0.0500732(53)$ $100.1(9.1)$ $186.5(5.6)$ $0.0500786(41)$ $123(10)$ $79.8(4.8)$
$H_{26}$ $F_{15}$ $0.0520673(53)$ $105.9(9.3)$ $208.5(5.5)$ $0.0520599(50)$ $110(10)$ $229.6(5.4)$
$H_{27}$ $F_{23}$ $0.0548976(91)$ $73(10)$ $41.1(8.4)$ $0.054928(11)$ $60(11)$ $228(10)$
$H_{28}$ $0.077011(20)$ $100(18)$ $227(10)$
$H_{29}$ $F_{13}$ $0.084545(23)$ $151(22)$ $184.2(9.0)$ $0.084530(21)$ $158(23)$ $357.8(8.3)$
$H_{30}$ $0.085006(30)$ $110(23)$ $68(12)$
$H_{31}$ $F_{12}$ $0.092850(34)$ $190(33)$ $26(10)$ $0.092665(33)$ $166(32)$ $0(11)$
$H_{32}$ $0.093638(44)$ $157(34)$ $85(13)$ $0.093750(29)$ $201(34)$ $78.1(9.7)$
$H_{33}$ $F_{3}$ $0.101556(19)$ $583(48)$ $67.6(5.1)$ $0.101567(15)$ $638(47)$ $141.5(4.3)$
$H_{34}$ $0.128854(66)$ $191(34)$ $237(11)$
$H_{35}$ $F_{7}$ $0.132814(40)$ $301(31)$ $226.6(6.3)$ $0.132831(30)$ $346(30)$ $338.9(5.1)$
$H_{36}$ $0.22254(17)$ $112(19)$ $126(10)$
$H_{37}$ $0.26299(26)$ $139(23)$ $216(10)$ $0.26367(29)$ $92(20)$ $8(12)$
$H_{38}$ $0.32205(42)$ $138(25)$ $114(11)$
$H_{39}$ $F_{5}$ $0.40112(28)$ $384(28)$ $358.4(4.8)$ $0.40099(17)$ $427(22)$ $349.3(3.1)$
$H_{40}$ $0.47653(65)$ $259(33)$ $81.9(7.8)$ $0.47670(67)$ $171(26)$ $127.1(8.3)$
$H_{41}$ $0.49322(64)$ $194(26)$ $140.6(7.5)$
$H_{42}$ $F_{2}$ $0.52602(31)$ $713(36)$ $233.2(3.1)$ $0.52575(21)$ $652(26)$ $316.7(2.3)$
$H_{43}$ $1.0516(26)$ $321(39)$ $188.2(7.1)$
NB (i) Uncertainties are given in parentheses after the estimation, in the unit of the latter’s last significant digit. This applies to all similar tables below.
NB (ii) Harmonics labelled with $F_{k}$ (col. 2) were cross-identified with those reported by von Essen et al. (2020); this label is the same as used in that paper.
In Fig. 4, top panels, we plot all the detected harmonics in the frequency – amplitude plane (with uncertainties). Harmonics that were cross-identified by their frequencies are additionally labelled with circles. We can see that in a given TESS sector just a few harmonics miss a counterpart in the another one. Such cases might indicate that the given harmonic is a false positive or vice versa, that the counterpart harmonic failed to overcome the $\rm FAP=5\%$ threshold level.
Figure 4: All the detected harmonics in the frequency – amplitude plane and their contributions to the total cumulative power.
In the bottom panels of Fig. 4 we plot the cumulative function of the harmonics power $A_{k}^{2}/2$ against the amplitude $A_{k}$. From these graphs we can see that $\sim 70\%$ of the total power (variance) of the compound harmonics signal is contained in just $4$ or $5$ components of largest amplitude ($A_{k}\gtrsim 600$ ppm), while the remaining $\sim 30\%$ of the total power is distributed among the rest ($A_{k}\lesssim 400$ ppm). The total power of the compound signal corresponds to $\sim 1200$ ppm for TESS18 and TESS58 both.
Table 2: Estimated noise parameters.
Parameter TESS18 TESS58
$\sigma_{0}$ [ppm] $260.4(5.3)$ $544.6(4.9)$
$\sigma_{1}$ [ppm] $525(47)$ $294(23)$
$\tau_{1}$ [d] $0.52(13)$ $0.230(98)$
$P^{\prime}_{1}$ [d] $1.65(23)$ $\infty(\gtrsim 1.7)$
$\sigma_{2}$ [ppm] $426(23)$ $470(22)$
$\tau_{2}$ [d] $0.182(26)$ $0.182(21)$
$P^{\prime}_{2}$ [d] $0.1085(15)$ $0.1084(13)$
Table 2 contains the best fitting NWN parameters of the model (2). The total noise variance corresponds to $\sim 700$ ppm and appears nearly the same for TESS18 and TESS58. Within it, the second NWN term appears nearly the same, although the first one reveals a significant change between the sectors, as well as the WN term does. But because of the stable total variance, this difference looks like a power exchange between the WN and the first NWN term over time. The total model variance becomes therefore $\sim 1400$ ppm for the both TESS sectors, and the variance of these data themselves matches this level closely, confirming that our models are consistent and close to being statistically exhaustive.
Figure 5: Power spectrum and correlation function of the compound model “harmonics+NWN”.
In Fig. 5 we show the power spectrum of the compound model “harmonics + NWN”, as well as its correlation function. The power spectrum is normalized so that it can be directly compared with the periodograms in Fig. 3. The NWN has a continuous power spectrum, but harmonics are expressed by delta functions there, so we plot them as vertical arrows starting from the NWN graph. In the correlation function plots we show separate graphs for the WN and NWN portions, and for the harmonics. Notice that correlation function of a single sinusoid in (1) is $\frac{A_{k}^{2}}{2}\cos 2\pi f_{k}\Delta t$.
4 Improving transit timing accuracy
4.1 Detrending and whitening the ITD data
Now let us consider how our out-of-transit photometric model of WASP-33 can be used to detrend the transit lightcurves. The most direct and robust way to achieve this is to fit the entire TESS dataset (OTD and ITD) with this model, augmented by the models of all the transits. However, this way it would require too slow nonlinear fits, and besides we have a goal in mind to employ our model with any other transit observations, so it is necessary to develop an approach that avoids using full TESS data directly.
Our model contains two components: the deterministic one $\mu(t)$ (harmonics) and the GP part (NWN) expressed by the correlation function $k(\Delta t)$. Regarding the deterministic part, its uncertainty can be computed based on the covariance matrix of fitted parameters. This uncertainty remains in the range $100$ ppm to $250$ ppm (for TESS18 and TESS58, resp.) Relatively to the average TESS-reported uncertainty for these data, $\sim 30000$ ppm, this appears below $1\%$. Therefore, possible uncertainty in $\mu(t)$ is surely negligible in comparison with the best expected photometric accuracy. So we can simply subtract our best fitting $\mu(t)$ from the ITD, as if it was a predefined function. We will refer this part of the ITD processing as detrending for shortness. However, notice that this differs from the detrending used in TESS data product (LC_DETREND), even though their practical effect may seem partly similar.
Regarding the NWN part of the model, its smallest $\tau_{k}$ from Table 2 are about $3$-$4$ h, which is similar to the exoplanet transit duration. This makes it tentative to construct some kind of a GP prediction model (based on the OTD subset) and then to subtract it from the ITD, just like $\mu(t)$. We undertook an attempt to compute an OTD-conditioned GP predictive model with parameters taken from Table 2. However, the practical predictability timescale appeared significantly smaller than $\tau_{k}$, namely below $1$ h. Along the transit duration the variance of this prediction remained mostly close to the total variance of the NWN (i.e. to the unconditional variance, as if OTD were not used). Therefore, such a predictive model would not be very helpful, but simultaneously this means that we would not loose significant information if we do not use the OTD directly when processing the ITD. This enabled us to apply a simplified approach with maximum-likelihood fitting of the ITD adopting a fixed covariance matrix (which is defined by the NWN parameters from Table 2).
Importantly, this approach is equivalent to the whitening in which the data are passed through a linear transform that renders them white. Formally, if $\textbf{{V}}=\mathop{\text{Var}}\nolimits\mn@boldsymbol{x}$ is the covariance matrix of the data ${x}$, and ${r}$ is the vector of their residuals then the modified (“whitened”) vector
${r}^{\prime}=\textbf{{V}}^{-\frac{1}{2}}\mn@boldsymbol{r}$ (5)
appears such that
$\mathop{\text{Var}}\nolimits\mn@boldsymbol{r}^{\prime}=\textbf{{I}},\quad|\mn@boldsymbol{r}^{\prime}|^{2}=\mn@boldsymbol{r}^{\rm T}\textbf{{V}}^{-1}\mn@boldsymbol{r}.$ (6)
The last quantity, |${r}^{\prime}|^{2}$, forms the only variable part of (3) when V is fixed. Therefore, with our GP-fitting algorithm it is unnecessary to apply whitening transform explicitly, but nonetheless we will refer this part of ITD processing as whitening. This should be conceptually the same as the whitening used in TESS data products (LC_WHITE), but in practice the result largely depends on the adopted V which was different.
Figure 6: Detrending and whitening for a few WASP-33 b transits. Notice that whitening normalizes the data to unit variance, changing their scale. Since this variance is always unit, the error bars for the whitened data are omitted for clarity. Models residuals are plotted in the bottom of each graph (with a constant subtracted). Same plots for all $38$ transits can be found in the online-only supplement.
Summarising, our ITD processing consists of two cleaning steps, detrending and whitening. In Fig. 6 we show how they work for a few sample transits. In these graphs we plot the TESS data before processing (green points), the sum-of-harmonics $\mu$ (orange curve), the detrended ITD (magenta points, shifted down) together with the transit model (orange curve), the corresponding residuals (shifted down yet more), and detrended and whitened ITD (blue points), also together with their model (which is also whitened) and residuals (shifted down). One can see that detrending can remove a large portion of the host star variability, however some residual nuisance variation still remains and should be removed by whitening.
4.2 Transit lightcurves fit
We applied an NWN algorithm for a self-consistent fit of all our $38$ transits, which was largely similar to the one presented in (Baluev et al., 2019). The flux model involved a cubic trend with fittable coefficients (individual per each transit), multiplied by the transit lightcurve model by Mandel & Agol (2002) with quadratic limb darkening. We omitted the outliers detection stage (since outliers were already removed in the DVT data), and adopted fittable (rather than fixed) limb darkening coefficients. The NWN was modelled, whenever appropriate, using the same GP technique and the same correlation models as described in Sect. 3.1.
Table 3: Best fitting WASP-33 system parameters.
Planet Star Limb darkening
$P$ [d] $1.219870527(70)$ $M_{\star}$ [$M_{\odot}$] $1.495(31)$∗ $A$ $0.212(34)$
$l$ [] $86.412(10)$ $R_{\star}$ [$R_{\odot}$] $1.5101(68)$ $B$ $0.097(54)$
$i$ [] $87.21(48)$ $\rho_{\star}$ [$\rho_{\odot}$] $0.4342(59)$ $A_{\rm I}$ $0.1883$
$r/R_{\star}$ $0.11078(25)$∗∗ $B_{\rm I}$ $0.2873$
$r$ [$R_{\rm Jup}$] $1.665(10)$ $A_{\rm II}$ $0.246(6)$
$a$ [AU] $0.02555(35)$ $B_{\rm II}$ $0.252(6)$
${}^{*\phantom{**}}$ Epoch $\rm BJD_{TDB}=2459000$.
${}^{**\phantom{*}}$ Taken from (Collier Cameron et al., 2010).
∗∗∗ This value is considerably affected by the third light correction, so its uncertainty may be unreliable.
Fitting the detrended and whitened ITD allowed us to obtain an improved set of WASP-33 system parameters. They are given in Table 3. This fit used the stellar mass estimate by Collier Cameron et al. (2010), but ignored their stellar radius value since we need to know the correlation between $M_{\star}$ and $R_{\star}$ to correctly include them both. Our estimate for the stellar density appeared somewhat smaller though (by $\sim 2.5\sigma$). This model assumes that transit timings strictly follow a linear ephemeris (no TTV).
In addition to the best fitting estimates of limb darkening coefficients, we also give in Table 3 their reference values (i) computed from the tables by Claret (2017) using $T_{\rm eff}$, $\log g$, and [Fe/H] by Collier Cameron et al. (2010) and (ii) computed by von Essen et al. (2020). These reference sets of estimates both differ from the fitted one, as well as from each other. This may indicate some issues regarding their reliability for not yet clear reasons.
4.3 Transit timings
On the final step we fitted the ITD once again, but assuming fittable transit timings. This was done for all three types of the ITD: raw, detrended and whitened detrended. The effect of our cleaning method is demonstrated in Fig. 7, where we plot the TTV residuals derived from the initial (raw) ITD, detrended ITD and detrended and whitened ITD. One can see that the TTV scatter is remarkably reduced after each step. The initial TTV residuals r.m.s. was $63$ s. After detrending it became $36$ s, while after the whitening it finally dropped to $18$ s. In total we have a rather dramatic r.m.s. improvement by a factor of $3.5$.
Figure 7: WASP-33 transit timing residuals before and after ITD cleaning. In the bottom pair of graphs the whitened detrended sets were replotted using larger ordinate scale.
Table 4: WASP-33 b transit timings derived from the TESS ITD.
Count Transit midtime Uncertainty Count Transit midtime Uncertainty
${\rm BJD}_{\rm TDB}-2450000$ [s] ${\rm BJD}_{\rm TDB}-2450000$ [s]
00 $8791.414035$ $22.7$ $895$ $9883.198699$ $17.7$
$1$ $8792.633942$ $14.0$ $896$ $9884.418604$ $18.0$
$2$ $8793.854405$ $15.3$ $897$ $9885.638319$ $21.2$
$3$ $8795.073867$ $14.4$ $898$ $9886.858225$ $20.3$
$4$ $8796.293688$ $13.0$ $899$ $9888.077433$ $22.2$
$5$ $8797.513648$ $13.8$ $900$ $9889.297612$ $20.2$
$6$ $8798.733613$ $14.7$ $901$ $9890.517565$ $18.1$
$7$ $8799.953544$ $13.0$ $902$ $9891.737567$ $19.2$
$10$ $8803.612739$ $13.0$ $903$ $9892.957195$ $19.2$
$11$ $8804.832991$ $14.8$ $904$ $9894.177081$ $19.2$
$12$ $8806.052820$ $13.8$ $905$ $9895.396834$ $20.9$
$13$ $8807.272818$ $14.9$ $907$ $9897.836626$ $18.3$
$14$ $8808.492436$ $15.2$ $908$ $9899.056929$ $18.3$
$15$ $8809.712360$ $15.9$ $909$ $9900.276757$ $19.0$
$16$ $8810.931993$ $13.2$ $910$ $9901.496700$ $18.5$
$17$ $8812.152098$ $13.7$ $911$ $9902.716293$ $19.5$
$18$ $8813.372200$ $14.7$ $912$ $9903.936345$ $15.9$
$913$ $9905.155751$ $16.8$
$914$ $9906.375920$ $18.0$
$915$ $9907.595623$ $19.3$
$916$ $9908.816113$ $19.1$
NB Machine-friendly version of this table, with uncertainties given in days, can be found in the online-only supplement.
The last best-accuracy set of transit timings is given in Table 4, while in Table 5 we provide some statistics regarding the fitted timings for all three types of the ITD, and separately for TESS sectors. One can see that only the full cleaning makes our fit-derived timing uncertainties to agree with their actual SSD. In other cases the SSD appears larger than formal uncertainties, indicating inaccuracy of the noise model. We also note that TESS data reveal remarkable homogeneity and stability over a single sector, which is indicated by very small values of uncertainty SSD (last column in Table 5), compared to the uncertainty mean value. There is some difference in the average uncertainty between sectors, but it corresponds to just about $1.6\sigma$, which is not enough for reliable conclusions.
Table 5: Statistics of the resulting transit timings and their uncertainties.
Data type Sector Mean timing residual [s] Residuals SSD [s] Mean timing uncertainty [s] Uncertainties SSD [s]
Raw 18 $-3.0$ $65.7$ $26.4$ $5.0$
58 $-4.4$ $64.3$ $29.5$ $4.3$
Detrended 18 $-3.8$ $33.4$ $13.5$ $1.8$
58 $-1.9$ $39.8$ $17.2$ $2.0$
Detrended + 18 $+0.2$ $15.8$ $14.7$ $2.2$
whitened 58 $-0.7$ $20.6$ $19.0$ $1.5$
Sample Standard Deviation.
We also investigated timing data from Table 4 for possible variations, but with a null result: we did not find any significant quadratic term in the TTV trend and we did not find any significant periodic TTV.
5 Discussion
The approach based on the explicit modelling of WASP-33 variability appears extremely successful in improving its exoplanet transit timings accuracy. In total, we obtained a dramatic TTV r.m.s. improvement by a factor of $3.5$. It is important that we must take into account the deterministic as well as non-deterministic (NWN) parts of this variability. Both of them contribute a similar amount in the total TTV error budget.
Our results still cannot be directly applied to observations from other sites than TESS, because the model parameters somewhat change on the timescale of a few years. Because of that we cannot use TESS-derived model with data obtained e.g. years prior to TESS. However, the frequencies of the harmonics and of the QP part of the NWN look stable, so it may be possible to apply our model to ground-based WASP-33 photometry assuming that only these frequencies are fixed and to refit harmonics amplitudes and phases, as well as NWN variances. But possible ways to reprocess old data using TESS-based models need more investigation.
Our results regarding the deterministic part of the WASP-33 variability are generally similar to von Essen et al. (2020). We found $43$ harmonics in total, although only $30$ of them are surely reliable because they were revealed in the both datasets, TESS18 and TESS58. von Essen et al. (2020) detected $29$ harmonics in TESS18, which is nearly the same number. However, there are a few individual differences. Based on our analysis we cannot confirm that their harmonics F21, F25, and F29 are statistically significant, even in a single TESS sector. On contrary, our harmonics H22, H32, H37, and H40 were revealed both in TESS18 and TESS58, but left undetected by von Essen et al. (2020). Among our harmonics that were revealed in TESS18, but not in TESS58, only one (H3 = F28) appeared in von Essen et al. (2020). Among harmonics that were detected in TESS58, but not in TESS18, none appeared in von Essen et al. (2020). This may render our $13$ “single-sector” harmonics less reliable, although this does not necessarily mean that all of them, or their majority, are false positives. Based on our FAP threshold of $5\%$, the expected number of false positives after $36$ or $37$ applications of the significance test should be just one or two (for each TESS sector). And these false positives are, most probably, among the “single-sector” harmonics.
Regarding the secondary eclipses of WASP-33 b, detected by von Essen et al. (2020), we cut them away from our OTD, so they did not affect our analysis and we did not consider them here. But two other planet-related effects reported by (von Essen et al., 2020), namely the orbital phase variation and the ellipsoidal variation, could affect our results, so we included them in the model $\mu$ from the beginning of our analysis. For the orbital phase variation our final amplitude estimate was $A_{\rm pv}=(110\pm 100)$ ppm and $(32\pm 44)$ ppm (resp. in TESS18 and TESS58), while von Essen et al. (2020) gave $A_{\rm pv}=(100.4\pm 13.1)$ ppm.
One can see that our uncertainty for $A_{\rm pv}$ is much larger and this requires investigation. We considered a number of possible explanations for such difference. First, our analysis used a reduced phase coverage because we cut away transits and secondary eclipses. In total, an $8$ hours portion of each orbital period was removed, which is only $27$ per cent of the full sector 18 data. This should leave about $90$ per cent of the $A_{\rm pv}$ variation, and clearly could not increase its uncertainty by a factor of $7$. Secondly, we tried to seek and cut away potentially unreliable parts of the data. In particular in the beginning of the TESS18 lightcurve the spacecraft passed through the Earth shadow that changed its temperature mode. Removing a small initial portion before the passage resulted in $A_{\rm pv}=(120\pm 100)$ ppm. Finally we considered the raw SAP data. We applied a simple detrending technique that was discussed in Sect. 2.2, again removed the segment before the Earth shadow passage, and then refitted our compound model “NWN+harmonics” using these data. We obtained $A_{\rm pv}=(130\pm 80)$ ppm after that. In fact we found the only way to reproduce $\sigma_{A_{\rm pv}}\sim 13$ ppm, which was to remove the NWN treatment entirely. By fitting the data as if the noise was purely white, we had $A_{\rm pv}=(130\pm 14)$ ppm.
This looked as if the issue is related to how the NWN was processed, and therefore we tried to reproduce von Essen et al. (2020) methods at least partly. We applied the $\beta$ method by Winn et al. (2008) to our OTD, which resulted in the maximum $\beta\simeq 11$ (achieved for the bin width about $1$ d). Given this value, we should scale our WN-derived uncertainty up from ±$\pm 14$ ppm to ±$\pm 150$ ppm. Therefore, if we used $\beta$ method in our study instead of GP fitting, we would obtain even larger uncertainty for $A_{\rm pv}$.
Detailed investigation of other parameters fitted by von Essen et al. (2020) revealed a yet another cautioning matter. Literally all amplitudes of the pulsation harmonics have exactly the same uncertainty of ±$\pm 10$ ppm (see their Table 2). However, such condition is impossible for correlated noise, because its magnitude varies over the frequency axis. In case of WASP-33 its power spectrum is peaked at low frequencies, so harmonics located in this range must have larger uncertainties than high-frequency ones, which are practically unaffected by the NWN. Such a behaviour is obvious in Table 1, where the uncertainties of $A_{k}$ increase by a factor of $\sim 5$ from top line to bottom.
The fact the all $\sigma_{A_{k}}$ appeared the same in (von Essen et al., 2020) might have two alternative explanations. The first one is due to the $\beta$ method which in fact does not offer a genuine NWN treatment: it simply scales all uncertainties up by the most conservative multiplier. In this case uncertainties of low-frequency harmonics would be more adequate, but those at high frequencies would be considerably overvalued. The opposite explanation is that high-frequency harmonics are adequate, while low-frequency ones are considerably less constrained than it follows from the stated uncertainties.
To determine which situation takes place here, let us assess the uncertainty of harmonics amplitudes, assuming that noise is white. There is a well-known approximation of this uncertainty (Vityazev, 1994, eqs 6.17-6.19) [^3] Formally, this approximation requires that the spectral window is small at all frequencies $f_{i}\pm f_{j}$, where $f_{i}$ being the $i$th harmonic frequency (this set includes $2f_{i}$, but excludes zero). The TESS time series are practically even, so the spectral window is always small in the domain of interest.:
$\sigma_{A}\simeq\sigma_{\rm noise}\sqrt{\frac{2}{N}}.$ (7)
From Fig. 10 by von Essen et al. (2020) the value of $\sigma_{\rm noise}$ follows to be about $700$ ppm, while $N$ was $14000$. Therefore, for a WN fit these data would imply $\sigma_{A}\simeq 8.4$ ppm. If NWN is being corrected with the $\beta$ method, a WN-based $\sigma_{A}$ should be multiplied by the factor $\beta=3.545$, as adopted by von Essen et al. (2020), resulting in $\sigma_{A}\sim 30$ ppm. This is basically a minimum realistic value for an analysis that honoured NWN. However, in actuality they give $\sigma_{A}=10$ ppm, close to the pure WN case.
Based on the investigation presented above we believe that by some hidden mistake, or because of a software bug, von Essen et al. (2020) uncertainties for the harmonics $F_{k}$ refer to the WN treatment instead of the NWN one. Most probably, the same holds true for the putative phase variation. This harmonic should be correlated with the transit and secondary eclipse parameters, and hence $\sigma_{A_{\rm pv}}$ may be even larger than $\sigma_{A_{k}}$. According to von Essen et al. (2020), $\sigma_{A_{\rm pv}}=13.1$ ppm is a bit larger than $\sigma_{A_{k}}=10$ ppm indeed, but not enough to be realistic for an NWN fit ($\sim 30$ ppm at least). If we assume that ±$\pm 13.1$ ppm refered actually to a WN model, the NWN uncertainty would be about ±$\pm 50$ ppm. We also believe that constant flux term $c_{0}$ has an unrealistically small uncertainty too, while other fit parameters are less obvious.
From our GP analysis we cannot confirm that there is a statistically significant evidence in favour of orbital phase variation in TESS data — not only because of a formally larger uncertainty we obtained, but also because in the second TESS sector this estimation appeared to differ from the first one.
For the ellipsoidal variation we obtained the amplitude of $(65\pm 41)$ ppm and $(97\pm 31)$ ppm, for the two TESS sectors. von Essen et al. (2020) gave $27.4$ ppm in the text, although we think this could be a typo, because from their Fig. 9 this amplitude follows to be $\sim 40$ ppm. The latter value is within at most two-sigma from our estimations, but still it is comparable to the uncertainty and cannot be reliably confirmed observationally. However, contrary to the phase variation, this amplitude is theoretically predicted and represents something that required a reduction, rather than observational estimation or confirmation.
Acknowledgements
We are grateful to the anonymous reviewers for their useful and constructive suggestions that helped to improve the manuscript.
Data availability
The data underlying this article are available in the MAST (Mikulski Archive for Space Telescopes), at https://archive.stsci.edu/, in the article and in its online supplementary material.

### v3

@@
 Planet Star Limb darkening
-$P$ [d] $1.219870527(70)$ $M_{\star}$ [$M_{\odot}$] $1.495(31)$∗ $A$ $0.212(34)$
+$P$ [d] $1.219870527(70)$ $M_{\star}^{\rm CC10}$ [$M_{\odot}$] $1.495(31)$∗ $A$ $0.212(34)$
 $l$ [] $86.412(10)$ $R_{\star}$ [$R_{\odot}$] $1.5101(68)$ $B$ $0.097(54)$
-$i$ [] $87.21(48)$ $\rho_{\star}$ [$\rho_{\odot}$] $0.4342(59)$ $A_{\rm I}$ $0.1883$
-$r/R_{\star}$ $0.11078(25)$∗∗ $B_{\rm I}$ $0.2873$
-$r$ [$R_{\rm Jup}$] $1.665(10)$ $A_{\rm II}$ $0.246(6)$
-$a$ [AU] $0.02555(35)$ $B_{\rm II}$ $0.252(6)$
-${}^{*\phantom{**}}$ Epoch $\rm BJD_{TDB}=2459000$.
-${}^{**\phantom{*}}$ Taken from (Collier Cameron et al., 2010).
-∗∗∗ This value is considerably affected by the third light correction, so its uncertainty may be unreliable.
-Fitting the detrended and whitened ITD allowed us to obtain an improved set of WASP-33 system parameters.
-They are given in Table 3.
-This fit used the stellar mass estimate by Collier Cameron et al.
-(2010), but ignored their stellar radius value since we need to know the correlation between $M_{\star}$ and $R_{\star}$ to correctly include them both.
-Our estimate for the stellar density appeared somewhat smaller though (by $\sim 2.5\sigma$).
+$i$ [] $87.21(48)$ $\rho_{\star}$ [$\rho_{\odot}$] $0.4342(59)$ $A^{\rm C17}$ $0.1883$∗∗
+$r/R_{\star}$ $0.11078(25)$∗∗∗ $M_{\star}^{\rm DR3}$ [$M_{\odot}$] $1.54(4)$∗∗ $B^{\rm C17}$ $0.2873$∗∗
+$r$ [$R_{\rm Jup}$] $1.665(10)$ $R_{\star}^{\rm DR3}$ [$R_{\odot}$] $1.59(3)$∗∗ $A^{\rm vE20}$ $0.246(6)$∗∗
+$a$ [AU] $0.02555(35)$ $B^{\rm vE20}$ $0.252(6)$∗∗
+${}^{*\phantom{***}}$ Epoch $\rm BJD_{TDB}=2459000$.
+${}^{**\phantom{**}}$ Taken from (Collier Cameron et al., 2010).
+${}^{***\phantom{*}}$ Not used in the fit and given for comparison only (see text).
+∗∗∗∗ This value is considerably affected by the third light correction, so its uncertainty may be unreliable.
+Fitting the detrended and whitened ITD allowed us to obtain an improved set of WASP-33 system parameters, given in Table 3.
 This model assumes that transit timings strictly follow a linear ephemeris (no TTV).
-In addition to the best fitting estimates of limb darkening coefficients, we also give in Table 3 their reference values (i) computed from the tables by Claret (2017) using $T_{\rm eff}$, $\log g$, and [Fe/H] by Collier Cameron et al.
-(2010) and (ii) computed by von Essen et al.
+The fit included the stellar mass estimate by Collier Cameron et al.
+(2010) which was based on their spectrum model, line-profile tomography, and ground-based transit data constraining stellar density.
+This stellar mass was accounted for through a penalty function added to the log-likelihood (see Galazutdinov et al., 2023).
+We ignored the accompanying stellar radius value since we need to know the correlation between $M_{\star}$ and $R_{\star}$ to correctly include them both.
+Our estimate for the stellar density appeared smaller by $\sim 2.5\sigma$ than that by Collier Cameron et al.
+(2010).
+Independent estimations of WASP-33 mass and radius are available in GAIA DR3 FLAME data table (Gaia Collaboration; Vallenari et al., 2023; Gaia Collaboration, 2024); they were added to Table 3 for comparison.
+Although these estimations use more accurate parallax, we did not use them in our fit, because contrary to Collier Cameron et al.
+(2010) they (i) are based on a rather low-resolution spectrum, (ii) did not include constraints from any transit observations of WASP-33 b.
+This resulted in a considerably larger uncertainties in GAIA $M_{\star}$, $R_{\star}$.
+We were also bothered by very low metallicity of $-0.56$ given in the same GAIA table and looking inconsistent with e.g.
+Collier Cameron et al.
+(2010).
+Yet another independent $M_{\star}$, $R_{\star}$ estimations provided by von Essen et al.
+(2020) have even larger uncertainties than GAIA ones.
+In addition to the best fitting estimates of limb darkening coefficients, we also give in Table 3 their reference values (C17) computed from the tables by Claret (2017) using $T_{\rm eff}$, $\log g$, and [Fe/H] by Collier Cameron et al.
+(2010) and (vE20) computed by von Essen et al.
 (2020).